OpenStudy (anonymous):

Is this true always sometimes or never? (x)(x)=x^2

OpenStudy (anonymous):

always

OpenStudy (anonymous):

it is a definition

OpenStudy (anonymous):

ALWAys because when you multiply both it would be ^

OpenStudy (anonymous):

x*x=x^1+1=x^2

OpenStudy (anonymous):

cause you multiplying the same x justlike 2^2 ur multiplying 2 two times

OpenStudy (anonymous):

what about (x)(x)=x

OpenStudy (anonymous):

tht would be x^2 cause you combining

OpenStudy (anonymous):

so its always true? (x)(x)=x?

OpenStudy (anonymous):

(x)(x)=x is not true because (x)(x) actually = x^2